A solenoid consists of 100 turns of a wire in a length of 0.200m. If the current in the wire is 3.00 A, what is the magnitude of the magnetic field along the axis of the solenoid? The solenoid has a cross-sectional area of 0.00200m2. If the current drops from 3.00 A to 0 in 1/60 s, what is the magnitude of the average induced emf in the solenoid?
Answer #1

L = 0.2 m , N =100 turns, i 1= 3A,i2 =0 , A = 0.002 m^2 , t = 1/60 s

Magnteic field by soeloid B =uoNi/L

B1 = (4*3.14*10^-6*100*3)/0.2

B1 =0.01884 T

B2 = 0

From faradays law induced emf V = -d(phi)/dt = A(B1-B2)/t

V = (0.002*0.01884)/(1/60)

V =0.00226 V

V =